# Ax-Prover: A Deep Reasoning Agentic Framework for Theorem Proving in Mathematics and Quantum Physics

### v2

@@ 3.2 MCP Tools
As described above, tool use is essential in our approach. We provide the LLM with access to tools via the MCP, a standard interface that lets LLM agents invoke external services in a uniform, controlled way [41]. We implement two categories of tools: Filesystem tools and Lean tools. Filesystem tools handle file operations such as read_file, write_file, and list_directory (see Appendix A.1). Lean tools allow Ax-Prover to perform a variety of actions crucial for theorem proving. We access these tools through the lean-lsp-mcp project [22], which provides a standardized interface to the Lean environment and ensures that Ax-Prover always operates on the latest version of Mathlib. This ensures that we can maintain compatibility for imports, theorem references, and proof construction without relying on the LLM’s knowledge of the version (or multiple incompatible versions) of Mathlib on which it was trained. The Lean tools we use fall into four main groups, summarized in Table 1.
@@
Note that lean_diagnostic_messages returns an error log containing scalars indicating the severity of the errors or messages the tool has returned: 0 if no error is found; 1 for erroneous/incorrect proofs; and 2 for a valid but incomplete proof, e.g. with sorry, warning, or linter error. A proof is considered correct and complete only if it compiles, no severity 1 errors are found, and no sorry or admit is present.
@@ 4 Datasets
While the application of LLMs to mathematical verification in Lean is evolving rapidly, the availability of comprehensive datasets remains limited. At present, only a few open-source datasets are available, with some of the most notable being MiniF2F [65], PutnamBench [55], and NuminaMath-LEAN [44]. These benchmarks include hard, high-level math problems from competitions such as the International Mathematical Olympiad (IMO) or the Putnam exam. Other datasets exist, but have clear limitations. For example, the Deepseek-Prover-V1 Train[20] includes 27k LLM-generated statements and proofs, but most of them are very simple, and can be solved in 2–3 line (on average) proofs. Lean Workbook [64] (57k) gathers LLM-generated formalizations of mathematical problems. While it reports a $93.5\%$ statement-level accuracy after filtering, subsequent analyses note that a nontrivial fraction of examples still suffer from semantic errors and hallucinations [38, 59], which limits its reliability.
Current benchmark datasets focus on mathematics and, even within this domain, are centered narrowly on high school and undergraduate-level competition problems. To enrich the current ecosystem and expand the coverage of Lean datasets, we create two datasets. The first one AbstractAlgebra (AA), is a Lean 4 dataset of problems drawn from standard abstract algebra textbooks. Unlike existing math benchmarks, which focus on undergraduate level competition-style puzzles, AA targets graduate or research-level mathematics, emphasizing deeper abstract concepts over lengthy step-by-step manipulations. The second dataset is QuantumTheorems (QT), which covers core topics in foundational quantum mechanics, with problems spanning from density matrices to scaling laws for quantum repeater networks. By bridging theoretical physics with formal verification methods, QT not only offers an unprecedented opportunity to test prover agents on quantum mechanics theorems, but it also represents a key step toward evaluating scientific reasoning models across any scientific discipline grounded in mathematics. In the section below, we provide more information about these two as well as other datasets we used for our experiments.
@@ 4.1 AbstractAlgebra
AbstractAlgebra is a curated dataset of 100 Lean problems extracted from the exercises in Dummit & Foote’s abstract algebra textbook [23]. The problems were extracted and formalized in Lean using an automatic pipeline (see details and examples in Appendix B.1). The dataset consists of two subsets: 50 easy problems from Chapter 1.1 and 50 intermediate problems from Chapters 1.2–2.5. The two categories thus reflect the increasing level of difficulty of the chapters in the book.
@@
In contrast, the AA dataset is aimed toward research-level mathematics, which involves deeper concepts with fewer reasoning steps per exercise. For instance, an AA problem may ask: Prove that every element $x=sr^{i}$ in the dihedral group $D_{n}$ has order 2. By presenting this kind of problems, AA fills the gap between AI-focused formalization efforts, which largely targets elementary mathematics, and the advanced topics studied by research mathematicians.
Finally, we stress that abstract algebra is foundational to much of mathematics, providing essential tools for research in number theory, geometry, topology, and beyond – indeed, 22 of the 32 primary mathematics categories on arXiv build upon it [1]. It also underpins important domains outside of mathematics, such as cryptography, physics, and chemistry. The broad foundational nature of abstract algebra underscores the importance of developing AI proof systems that perform well on problems in this domain, as this has the potential to accelerate progress across many scientific fields.
@@ 4.2 QuantumTheorems
QuantumTheorems includes 134 problems spanning core areas of quantum theory. These problems introduce unique challenges, as they require integrating finite-dimensional linear algebra, complex analysis, and matrix theory with quantum principles such as unitarity, Hermiticity, and measurement postulates. This domain-specific knowledge is absent from existing datasets of formal proofs, making QT a valuable benchmark for testing and advancing formal reasoning in physics. QT was generated through an iterative human-in-the-loop process, combining automated proof synthesis with expert curation (see Appendix B.2 for more details and examples).
We generated problems at two levels of difficulty: basic problems are short (require on average 1–10 line proofs to solve) and often solvable with standard automation tactics (simp, linarith), e.g., a post-measurement state is an eigenstate of the measurement projector. Intermediate level problems (requiring 10–50 line proofs to solve) are solvable with systematic case analysis and orchestration of rewrite rules, such as proving simultaneous diagonalization of commuting observables.
QT represents a first step toward computer-verified quantum mechanics, addressing the challenge of ensuring correctness in quantum information protocols and algorithms. The dataset has practical importance beyond research: as quantum technologies grow more complex, errors in proofs or hidden assumptions can have serious consequences. For instance, a recent bug in a proof claiming to break lattice-based cryptography – only identified weeks later by experts – illustrates the risks of unchecked reasoning in high-stakes domains [50, 15]. QT provides a first-of-its-kind resource for developing tools which can help detect these mistakes earlier.
@@ 4.3 NuminaMath-LEAN
NuminaMath-LEAN [44] is a very recent (August 2025) large-scale collection of approximately 104,000 competition-level mathematics problems formalized in Lean 4. The dataset is created by the same research group that developed the Kimina-Prover. They derived NuminaMath-LEAN from NuminaMath 1.5 [34], with problems drawn from prestigious contests such as the International Mathematical Olympiad (IMO) and the United States of America Mathematical Olympiad (USAMO).
Each problem includes a formal statement in Lean 4, written either by a human annotator (19.3% of the problems) or by an autoformalizer model (80.7%) [44]. Out of the total problems, 25% were correctly proved by Kimina-Prover during its reinforcement learning (RL) training phase (Solved-K), 11% were proved by humans (Solved-H), while the remaining 64% do not have any proof (Unsolved) [58, 34, 44]. We analyzed problems across the three groups and observed a clear difficulty gradient: Solved-K << Solved-H << Unsolved. This ordering aligns with the fact that Solved-H and Unsolved problems could not be handled by Kimina-Prover, providing an implicit measure of difficulty. The fact that Solved-H proofs are on average longer than those in Solved-K (155 vs. 98 lines) also offers quantitative evidence consistent with our qualitative assessment. For our experiments, we randomly sampled 300 problems – 100 each from Solved-K, Solved-H, and Unsolved – to create a balanced, representative, and more budget-friendly benchmark.
@@ 4.4 PutnamBench
PutnamBench [55] is a multi-language benchmark designed to evaluate the ability of neural theorem provers to solve undergraduate-level competition mathematics problems. It includes formalizations of problems from the William Lowell Putnam Mathematical Competition (1962–2024) across three major proof assistants – Lean, Isabelle, and Rocq. The Lean subset contains 660 formalized problems, which we focus on in this work. The problems span a broad range of undergraduate topics, including Algebra, Analysis, Number Theory, Geometry, Linear Algebra, Combinatorics, Abstract Algebra, Probability, and Set Theory.
Unlike MiniF2F, which is now saturated, PutnamBench remains a challenging benchmark for most provers. Moreover, since it is widely adopted by many models, it serves as a high-value testbed for evaluating our approach against the best theorem-proving models currently available.
@@ 5 Experiments
In this section, we provide details about the experimental setup we implemented (Section 5.1) and the results (5.2), followed by an analysis of tool usage (5.3) and the challenges and costs of model deployment (5.4).
@@ 5.1 Experimental Setup
In the tests with New Benchmarks, we evaluated the performance of the Ax-Prover against three strong baselines:
@@
We evaluated all models using pass@1: While this idea is in sharp contrast with previous studies assessing pass with very high values (see, e.g., [49]), we believe it mirrors real-world usage, where researchers are constrained by time and budget limits, and cannot run a prover multiple independent times in the hope that one succeeds. For transparency and reproducibility, we note that while pass@1, for all the baselines, means trying to formalize the entire proof in a single shot, for Ax-Prover it means performing a sequence of steps (i.e., API calls) in which reasoning and tool calls are interleaved in a singular attempt to build the final proof–i.e., with no parallelization. For these experiments, we powered Ax-Prover using Claude Sonnet 4 [4]. Furthermore, to stay within budget, we capped Ax-Prover API calls at 200 and set a 25-minute timeout. For all models, we computed the final results by running an external Lean compiler on the generated files, and considered a proof correct if it compiled and included no sorry.
@@
The second benchmark group, which includes PutnamBench only, aims to evaluate Ax-Prover on one of the most challenging public benchmarks and compare its performance against existing state-of-the-art provers. Accordingly, we did not run baselines and instead directly compared our results with those reported on the official leaderboard [54]. For this test, we powered Ax-Prover with Sonnet-4.5, removed the 25-minute timeout, and increased the max number of API calls at 400, while still running it with pass@1, as defined above.
5.2 Results
Dataset Subset Ax-Prover Sonnet DS-Prover Kimina
NuminaMath solved-K 81% 7% 48% 100%
solved-H 47% 8% 14% 0%
unsolved 26% 1% 18% 0%
total 51% 5% 28% 31%
AbstractAlgebra easy 72% 10% 26% 12%
intermediate 56% 6% 22% 14%
total 64% 8% 24% 13%
QuantumTheorems easy 100% 54% 88% 72%
intermediate 92% 18% 48% 34%
total 96% 40% 61% 57%
Table 2: Models’ performance on NuminaMath, AA, and QT. The results on NuminaMath for Kimina are reported from  [44], and where obtained during its RL training phase with, on average, pass@68.
New Benchmarks
We report the results for this group in Table 2. On the Numina dataset, Ax-Prover scored 51% accuracy, outperforming DS-Prover (28%) and Kimina (31%) by a similar margin, while Sonnet only got 5% accuracy. Particularly relevant is the performance of Ax-Prover on Solved-H, where it solves almost half of the problems, and on Unsolved (26%). Notably, due to autoformalization (see Section 4.3), some theorems are ill-posed: during testing, Ax-Prover spotted them, and reported the error (see Appendix C).
On AA the gap in performance is striking, with Ax-Prover (64%) outperforming DS-Prover by 40%, while both Kimina (13%) and Sonnet get a very poor performance (8%). We suggest this is because the AA dataset is largely out-of-distribution for DS-Prover and Kimina. In fact, these models are trained primarily on Mathlib, which covers only a minimal subset of abstract algebra, or on undergraduate competition-level math problems, which are qualitative differently from those in AA (See Section 4.1).
On the QT dataset, Ax-Prover achieves perfect performance on the easy split and 92% accuracy on the intermediate split, yielding 96% accuracy overall. This represents a substantial gap compared to DS-Prover (61%) and Kimina (57%), with Sonnet falling well behind at 40%. Also in this case, we suggest that the performance gap stems from our approach’s flexibility to adapt across scientific domains, in contrast to the over-specialization of specialized models. To showcase the differences between the models, let’s consider the proofs that quantum observables are Hermitian matrices (full proofs available in Appendix D.1). DS-Prover misused the Hermitian field, misunderstanding its type, while Sonnet made a more sophisticated effort but encountered a rewrite pattern mismatch, which highlights its difficulties in managing Lean environment. In contrast, Ax-Prover succeeded through a systematic approach, explicitly applying the Hermitian property to diagonal elements, using the definition of conjugate transpose, and connecting it to the fact that a complex number equal to its conjugate is real. The case highlights that successful formal theorem proving requires careful, step-by-step reasoning, a solid grasp of type theory, and familiarity with library theorems – demonstrating that clarity and correctness outweigh clever shortcuts in formal verification.
PutnamBench
Table 3 reports the results for the top 10 scorers on PutnamBench. [^3] See the full leaderboard at [54]. In the “Compute” column, pass@ indicates the number of independent attempts to solve a proof. avg. pass@ is used for Hilbert, an agentic framework that parallelizes reasoning and verification at different levels [57]. The exact definition of this metric is unclear; our best assumption is that it reflects the average number of calls to Hilbert’s sub-agents. Similarly, medium refers to a specific test setup for Seed-Prover, in which the model is evaluated with parallelized refinement processes [13]. On this benchmark, Ax-Prover achieves 14% accuracy, placing third in the table. While this result is lower than the top scorers, it is important to note that it was obtained at a fraction of the cost of the top-performing models (see the “Compute” column).
Model Accuracy Compute
Hilbert 72% [462] avg. pass@1840
Seed-Prover 51% [329] medium
Ax-Prover 14% [92] pass@1
Goedel-Prover-V2 13% [86] pass@184
DeepSeek-Prover-V2 7% [47] pass@1024
DSP+ 4% [23] pass@128
Bourbaki 2% [14] pass@512
Kimina-Prover-7B-Distill 2% [10] pass@192
Self-play Theorem Prover 1% [8] pass@3200
Goedel-Prover-SFT 1% [7] pass@512
Gemini-2.5-Pro 0.5% [3] pass@1
GPT-4o 0.2% [1] pass@10
Claude-3.7-Sonnet 0% [0] pass@1
Table 3: Accuracy results on PutnamBench (% and absolute number of solved problems). $\dagger$ Remember that for Ax-Prover, pass@1 is made of multiple steps, see Section 5.1.
Overall, the results in this section indicate that Ax-Prover delivers strong performance across the board, ranking among the top models in mathematics and outperforming others in physics. Also, they highlight two key limitations of current approaches: specialized provers fail to generalize beyond their training domains, while general-purpose LLMs, while creative, cannot produce rigorous Lean proof. The fact that Ax-Prover nearly doubles the performance of the standalone LLM using the same model (Sonnet) demonstrates that combining agentic reasoning with Lean tool integration is essential for robust theorem proving across domains. We examine this aspect in more detail in the next section.
5.3 Analysis of Tool Usage
To measure the impact of tool usage on our approach, we analyzed the tool calls done by the Prover over the 100 problems we tested on the challenging NuminaMath Unsolved subset. We found that the Prover made an average of 100.76 tool calls per run. Tool usage is highly reliable, with success rates above 99%. [^4] The only exception is search in Filesystem tools, with 80%. However, this results from the Prover searching for files that do not exists.
Table 4 reports the 10 most frequently used tools. At the top is edit_file, as the Prover updates the Lean file at each step. lean_diagnostic_messages follows, reflecting explicit instructions to verify each proof step. lean_goal exposes the current proof state, while lean_loogle and lean_leansearch enable the Prover to search for relevant theorems in the library. Importantly, these tools are used autonomously, without additional guidance. Collectively, these statistics illustrate how Ax-Prover leverages a tight feedback loop of editing, goal inspection, search, and diagnostics.
Tool Calls
edit_file 36.79
lean_diagnostic_messages 30.73
lean_goal 8.17
lean_loogle 5.88
lean_leansearch 4.32
file_contents 3.00
write_file 2.71
read_text_file 2.24
lean_run_code 2.05
lean_hover_info 1.76
Table 4: Tool usage statistics.
Our assumption is that tool usage enhances proof quality by allowing Ax-Prover to use less common tactics. To test this, we analyzed the unique tactics used in the proofs generated by Ax-Prover, Kimina, and DeepSeek, under the hypothesis that a larger set of tactics reflects greater creativity (see the full list of tactics per model in Table 5). The three models share 28 tactics, but Ax-Prover uses 9 tactics not employed by DS-Prover or Kimina, whereas these models combine to use only three tactics absent in Ax-Prover. This finding supports our hypothesis that integrating frontier LLMs with Lean tools enhances creative exploration in proof construction.
5.4 Deployment Analysis
Besides performance, deployment complexity is critical when using AI models in real-world scenarios. Here we compare prover systems in this respect. DS-Prover and Kimina require GPU-accelerated, high-spec machines and are not available through model as a service (MaaS) providers. [^5] For instance, while DeepSeek-Prover-V2-671B was previously hosted by Novita [21], this endpoint now redirects to the general DeepSeek-V3 model We hosted DeepSeek and Kimina on Google Cloud: DeepSeek on an A3 Ultra VM with eight H200-141GB GPUs, and Kimina on an A2 High GPU VM with eight A100-40GB GPUs. Deployment is burdensome and demands MLOps expertise – users must match hardware specs, configure distributed runtimes, debug serving issues, and contend with scarce GPU availability, since cloud providers enforce strict quotas and long queues for H100/H200 GPUs. This hinders reproducibility even for well-funded teams. In contrast, Ax-Prover relies only on API calls, requiring no infrastructure beyond basic client access, and it can be executed locally on a client machine or remotely in a lightweight container.
On monetary costs, running DS-Prover and Kimina on 1000 datapoints cost approximately \$300 and \$2000, respectively, while Ax-Prover cost about \$4000. At first glance, our approach appears more expensive but only because we evaluate specialized models with pass@1. Had we followed the common practice of running them with higher pass@n values, the cost of these specialized models would have far exceeded ours. Furthermore, consider that on PutnamBench, DS-Prover was run with a pass@1024, thus leveraging way more computational resources, and only got 47 correct theorems, while Ax-Prover got 92. Moreover, general-purpose LLMs are on a rapid trajectory of improvement: each new generation delivers stronger reasoning at lower cost, suggesting that the relative efficiency of Ax-Prover will only increase over time.
The deployment and cost barriers of specialized models also help explain why they have not achieved widespread use beyond benchmark settings such as IMO-style mathematics problems. For most researchers, the need to manage specialized hardware, navigate GPU quotas, and bear high costs makes these systems effectively unusable in practice. Ax-Prover is more accessible to researchers not only because it eliminates these barriers, but also because it was explicitly designed to act as a supportive assistant, as we show in the next section.
6 Use Case: Researcher-Friendly Verification
Ax-Prover is able to engage in productive collaborations with human researchers by verifying intermediate proof steps, providing precise feedback, and guiding the overall direction of the proof. This capability differentiates our framework from existing specialized provers, which generally attempt to complete proofs in a single pass without any interaction with their external environment. Thus, Ax-Prover should be viewed not merely as a backend system but as an active collaborator in mathematical and scientific reasoning. In this Section, we present a concrete demonstration of Ax-Prover’s capabilities as an assistant.
One of the authors of this paper is an expert mathematician who collaborated with Ax-Prover to prove a recent cryptography result, A New Algorithm for Computing Branch Number of Non-Singular Matrices over Finite Fields [40] (see the full case study in Appendix F). The task involved two main steps, namely: formalizing the paper’s statements in Lean, and verifying the main claim. The mathematician jointly worked with Ax-Prover to structure the proof, validate lemmas, and complete the verification. Ax-Prover supported the whole process by checking intermediate lemmas and guiding proof strategies. In what came as a surprise, it also revealed an error in the original approach. This shows that Ax-Prover can not only reproduce known mathematical results at the cutting edge of research, but also advance the state of knowledge by providing formal verification of correct results and expose incorrect ones. Notably, the whole process lasted two working days and was carried out locally on the laptop of the mathematician. The outcome of the effort is a proof of over 2000 lines of Lean code. [^6] The resulting Lean blueprint can be made available upon request.
To better understand the value of Ax-Prover’s contribution to the proof defined above, consider the formalization of the Prime Number Theorem. Terence Tao and Alex Kontorovich initiated the Lean translation, but the project was only completed weeks later by Math, Inc.’s Gauss agent running on Morph.AI’s Infinibranch cluster [39]. While ultimately successful, this represented a massive engineering effort spanning several weeks for a well-known proof of comparable difficulty to our cryptography case study, which, by contrast, required only two days, a single mathematician, and a laptop.
This comparison underscores the power of Ax-Prover in enabling fast and efficient verification of research-level proofs of new results. Besides its ability to perform end-to-end tasks, Ax-prover is able to act as a collaborative teammate, by exposing intermediate reasoning, suggesting next steps, and accepting and providing guidance.

### v3

@@ 3.2 MCP Tools
As described above, tool use is essential in our approach. We provide the LLM with access to tools via the MCP, a standard interface that lets LLM agents invoke external services in a uniform, controlled way [modelcontextprotocol2024]. We implement two categories of tools: Filesystem tools and Lean tools. Filesystem tools handle file operations such as read_file, write_file, and list_directory (see Appendix A.1). Lean tools allow Ax-Prover to perform a variety of actions which are crucial for theorem proving. We give Ax-Prover access to these tools through the lean-lsp-mcp project [lean-mcp], which provides a standardized interface to the Lean environment. Access to these tools allows Ax-Prover to search within the local library, diagnose errors and warnings as they arise, observe the Lean context at any point in a proof, and query external search engines. Notably, external search engines provide Ax-Prover with more up-to-date information about Mathlib than what exists within an LLM’s parametric knowledge: Loogle searches for declarations within the newest version of Mathlib, and Leansearch is based on a past but recent version of Mathlib. Since Mathlib is a rapidly evolving library, this capability of Ax-Prover ensures compatibility for imports, theorem references, and proof constructions without relying on the LLM’s knowledge of the version (or multiple versions) of Mathlib on which it was trained. The Lean tools we use fall into four main groups, summarized in Table 1.
@@
Note that lean_diagnostic_messages returns a number: 0 for a successfully compiled proof with no errors or warnings; 1 for an explicit compilation error in the proof; and 2 for a successfully compiled proof with warning messages, e.g. the proof is incomplete (containing sorry), or the code style does not pass the linter test. A proof is considered correct and complete only if a code 0 is returned, or if a level 2 code is returned that does not contain sorry.
@@ 4 Datasets
While the application of LLMs to mathematical verification in Lean is evolving rapidly, the availability of comprehensive datasets remains limited. At present, only a few open-source datasets are available, with some of the most notable being MiniF2F [minif2f], PutnamBench [tsoukalas2024putnambench], and NuminaMath-LEAN [numinamath-lean]. These benchmarks include hard, high-level math problems from competitions such as the International Mathematical Olympiad (IMO) or the Putnam exam. Other datasets exist, but have clear limitations. For example, Deepseek-Prover-V1 Train [deepseek-prover-v1-dataset] includes 27k LLM-generated statements and proofs, but most of them are very simple and can be solved in 2–3 lines on average. Lean Workbook [lean-workbook] (57k) gathers LLM-generated formalizations of mathematical problems. While it reports a $93.5\%$ statement-level accuracy after filtering, subsequent analyses note that a nontrivial fraction of examples still suffer from semantic errors and hallucinations [lu2024process, autoformalization-survey], which limits its reliability.
Notably, current valuable benchmarks focus on mathematics and, even within this domain, are centered narrowly on high school and undergraduate-level competition problems. To enrich the ecosystem and expand the coverage of Lean datasets, we create two datasets. The first one AbstractAlgebra (AA), is a Lean 4 dataset of problems drawn from standard abstract algebra textbooks. Unlike existing math benchmarks, which focus on undergraduate level competition-style puzzles, AA targets graduate or research-level mathematics, emphasizing deeper abstract concepts over lengthy step-by-step manipulations. The second dataset is QuantumTheorems (QT), which covers core topics in foundational quantum mechanics, with problems spanning from density matrices to scaling laws for quantum repeater networks. By bridging theoretical physics with formal verification methods, QT not only offers an unprecedented opportunity to test prover agents on quantum mechanics theorems, but also represents a key step toward evaluating scientific reasoning models across any scientific discipline grounded in mathematics. In the section below, we provide more information about these two datasets as well as the other datasets we used for our experiments.
@@ 4.1 AbstractAlgebra
AbstractAlgebra is a curated dataset of 100 Lean problems extracted from the exercises in Dummit & Foote’s abstract algebra textbook [dummit2004abstract]. The problems were extracted and formalized in Lean using an automated pipeline (see details and examples in Appendix B.1). The dataset consists of two subsets: 50 easy problems from Chapter 1.1 and 50 intermediate problems from Chapters 1.2–2.5. The two categories thus reflect the increasing level of difficulty of the chapters in the book.
@@
In contrast, the AA dataset is aimed toward research-level mathematics, which involves deeper concepts with fewer reasoning steps per exercise. For instance, an AA problem may ask: Prove that every element $x=sr^{i}$ in the dihedral group $D_{n}$ has order 2. By presenting these kinds of problems, AA fills the gap between AI-focused formalization efforts, which largely target elementary mathematics, and the advanced topics studied by research mathematicians.
Finally, we stress that abstract algebra is foundational to much of mathematics, providing essential tools for research in number theory, geometry, topology, and beyond – indeed, 22 of the 32 primary mathematics categories on arXiv build upon it [arxivMathArchive]. It also underpins important domains outside of mathematics, such as cryptography, physics, and chemistry. The broad foundational nature of abstract algebra underscores the importance of developing AI proof systems that perform well on problems in this domain, as this has the potential to accelerate progress across many scientific fields.
@@ 4.2 QuantumTheorems
QuantumTheorems includes 134 problems spanning core areas of quantum theory. These problems introduce unique challenges, as they require integrating finite-dimensional linear algebra, complex analysis, and matrix theory with quantum principles such as unitarity, hermiticity, and measurement postulates. This domain-specific knowledge is absent from existing lean datasets, making QT a valuable benchmark for testing and advancing formal reasoning in physics. QT was generated through an iterative human-in-the-loop process, combining automated proof synthesis with expert curation (see Appendix B.2 for more details and examples).
We generated problems at two levels of difficulty: basic problems are short (proof requires 1-10 lines of Lean code) and often solvable with standard automation tactics (simp, linarith), e.g., proving that a post-measurement state is an eigenstate of the measurement projector. Intermediate level problems (proof requires 10-50 lines of Lean code) are solvable with systematic case analysis and orchestration of rewrite rules, such as proving simultaneous diagonalization of commuting observables.
QT represents a first step toward computer-verified quantum mechanics, addressing the challenge of ensuring correctness in quantum information protocols and algorithms. The dataset has practical importance beyond research: as quantum technologies grow more complex, errors in proofs or hidden assumptions can have serious consequences. For instance, a recent bug in a proof claiming to break lattice-based cryptography – only identified weeks later by experts – illustrates the risks of unchecked reasoning in high-stakes domains [rousseau2024bug, chen2024lattice]. QT provides a first-of-its-kind resource for developing tools which can help detect these mistakes earlier.
@@ 4.3 NuminaMath-LEAN
NuminaMath-LEAN [numinamath-lean] is a very recent (August 2025) large-scale collection of approximately 104,000 competition-level mathematics problems formalized in Lean. The dataset is created by the same research group that developed the Kimina-Prover [kimina-prover]. They derived NuminaMath-LEAN from NuminaMath 1.5 [numina_math_datasets], with problems drawn from prestigious contests such as the International Mathematical Olympiad (IMO) and the United States of America Mathematical Olympiad (USAMO).
Each problem includes a formal statement in Lean, written either by a human annotator (19.3% of the problems) or by an autoformalizer model (80.7%) [numinamath-lean]. Out of the total problems, 25% were correctly proved by Kimina-Prover during its reinforcement learning (RL) training phase (Solved-K), 11% were proved by humans (Solved-H), while the remaining 64% do not have any proof (Unsolved) [kimina-prover, numina_math_datasets, numinamath-lean]. We analyzed problems across the three groups and observed a clear difficulty gradient: Solved-K << Solved-H << Unsolved. This ordering aligns with the fact that Solved-H and Unsolved problems could not be solved by Kimina-Prover, providing an implicit measure of difficulty. The fact that Solved-H proofs are on average longer than those in Solved-K (155 vs. 98 lines) also offers quantitative evidence consistent with our qualitative assessment. For our experiments, we randomly sampled 300 problems – 100 each from Solved-K, Solved-H, and Unsolved – to create a balanced, representative, and more budget-friendly benchmark.
@@ 4.4 PutnamBench
PutnamBench [tsoukalas2024putnambench] is a multi-language benchmark designed to evaluate the ability of neural theorem provers to solve undergraduate-level competition mathematics problems. It includes formalizations of problems from the William Lowell Putnam Mathematical Competition (1962–2024) across three major proof assistants – Lean, Isabelle, and Rocq. The Lean subset contains 660 formalized problems, which are the ones we focus on in this work. The problems require the clever application of a wide range of undergraduate topics, including abstract algebra [^4] The abstract algebra problems in PutnamBench are of a different nature than the ones in our AbstractAlgebra dataset, which focuses more on detailed facts about textbook level concepts requiring few steps of reasoning., analysis, number theory, geometry, linear algebra, combinatorics, probability, and set theory. For each year, the Putnam competition includes two sessions of six problems each, labeled A1–A6 and B1–B6. It is generally accepted that, within each session, the problems increase in difficulty from 1 to 6. Unlike MiniF2F, which is now saturated (see, e.g., [deepseek-prover-v2]), PutnamBench remains a challenging benchmark for most provers. Moreover, since it is widely adopted by many models, it serves as a high-value testbed for evaluating our approach against the best theorem-proving models currently available.
@@ 5 Experiments
In this section, we provide details about the experimental setup we implemented (Section 5.1) and our results (LABEL:subsec:results), followed by an analysis of tool usage (LABEL:subsec:analysis_tools) and the challenges and costs of model deployment (LABEL:subsec:deployment_analysis).
@@ 5.1 Experimental Setup
In the tests with New Benchmarks we evaluated the performance of the Ax-Prover against three strong baselines:
@@
We evaluated all models using pass@1: while this idea is in sharp contrast with previous studies assessing provers with pass@ very high values (see, e.g., [deepseek-prover-v2]), we believe it mirrors real-world usage, where researchers are constrained by time and budget limits, and cannot run a prover multiple independent times in the hope that one succeeds. For transparency and reproducibility, we note that while pass@1, for all the baselines, means trying to formalize the entire proof in a single shot, for Ax-Prover it means performing a sequence of steps (i.e., API calls) in which reasoning and tool calls are interleaved in a singular attempt to build the final proof, i.e. without multiple independent attempts (see Section 3.1.2). For these experiments, we powered Ax-Prover using Claude Sonnet 4 [anthropic2025claude4]. Furthermore, to stay within budget, we capped Ax-Prover API calls at 200 and set a 25-minute timeout. For all models, we computed the final results by running an external Lean compiler on the generated files, and considered a proof correct if it compiled and included no sorry.
@@
The second benchmark group, which includes PutnamBench only, aims to evaluate Ax-Prover on one of the most challenging public benchmarks and compare its performance against existing state-of-the-art provers. Accordingly, we did not run baselines and instead directly compared our results with those reported on the official leaderboard [putnambench_leaderboard]. For this test, we powered Ax-Prover with Sonnet 4.5, removed the 25-minute timeout, and increased the max number of API calls to 400, while still running Ax-Prover with pass@1, as defined above.